# Monitoring Risks in Test-Time Adaptation

### v1

@@ 1 Introduction
Test-time adaptation (TTA) [44] has proven to be a powerful paradigm for prolonging the life of a model subjected to distribution shift. TTA methods adapt model parameters online, using only test batches of features and no labels. By leveraging unsupervised objectives such as test-time entropy [42] or pseudo-label losses [41], these methods effectively ‘fine-tune’ model parameters on unlabeled test data. Despite their stark potential to retain high performance under a variety of distribution shifts [49], TTA methods can suffer performance drops under severe shifts or after prolonged adaptation. The TTA literature has documented a range of persistent failure cases in which the model collapses entirely, resulting in near-zero accuracy [31]. Alarmingly, these failures often occur silently and prohibit TTA methods from safe deployment in practice.
Timely detection of performance degradations—whether due to harmful distribution shifts or adaptation collapse—is thus crucial for safe deployment. At the same time, however, falsely flagging that a model should be taken offline and retrained can incur significant, avoidable costs given the size of modern predictive models. Recently, sequential testing has emerged as a promising statistical framework for monitoring model performance over time [30]. When a predefined risk (or error) threshold is exceeded, the monitoring tool triggers an alarm. By employing time-uniform confidence sequences [13], such tools provide rigorous guarantees on the false alarm rate, thereby minimizing unnecessary retraining with high probability.
However, existing sequential risk monitors either require access to ground truth labels in production [30, 51] or do not account for model adaptation [1]. In this paper, we extend sequential testing to the challenging setting of TTA. This enables us to track the risk of a continuously evolving model without ever observing test labels. Our main contributions are as follows:
@@
  • In § 3.2 and § A.1, we extend prior work on unsupervised risk monitoring [1] to enable effective tracking of risks most commonly used in TTA, such as classification error.
@@ 2 Preliminaries
When averaged across samples, this squared loss corresponds to the Brier score, a strictly proper scoring rule [9]. Hence it captures not only the classifier’s error but also its calibration. Due to this connection, we refer to this second loss as the ‘Brier loss’ for short.
@@
To track how well a deployed model is performing, Podkopaev and Ramdas [30] propose a risk monitoring framework based on sequential testing [32]. The performance of a model $p$, in the presence of a labeled test stream, is tracked using the following sequential test:
@@
To give the test anytime-valid properties (e.g. arbitrary stopping and restarting), Podkopaev and Ramdas [30] rely on confidence sequences, which extend traditional confidence intervals to the sequential setting and offer time-uniform coverage guarantees [5, 13]. A sequence of model losses on test data is used to construct an anytime-valid lower bound $L_{t}$ on the true running test risk $\bar{R}_{t}$:
@@
$\hat{U}=\hat{R}_{0}(p;\mathcal{D}_{\text{cal}})+w_{0}\>,\quad\hat{L}_{t}=\frac{1}{t}\sum_{k=1}^{t}\hat{R}_{k}(p;\mathcal{D}_{{\bm{x}}y}^{k})-w_{t}\>,$
@@
Figure 1: Alarm $\Phi_{t}$ is raised at $t_{\text{min}}$ as the lower bound $\hat{L}_{t}$ on the running test risk $\hat{\bar{R}}_{t}$ exceeds the upper bound $\hat{U}_{0}$ on the source risk $\hat{R}_{0}$.
which ensures that performance degradations are not incorrectly detected, thereby avoiding unnecessary (and potentially costly) interventions on the model. Notably, this guarantee requires only that the loss function is bounded $\ell\in[a,b]$. This minimal assumption makes the method broadly applicable, as it imposes no constraints on the data distributions, the predictive model, or the nature of distribution shift (beyond independence). To maintain minimal assumptions, it is necessary to rely on a conjugate-mixture empirical Bernstein bound [13] when constructing a lower confidence sequence for the test risk $L_{t}$ (see § B.1 for more details).
@@
In test-time adaptation [42], the model parameters $\theta$ are updated online as the model observes batches of unlabeled test features. Specifically, given a sequence of unlabeled test batches $\mathcal{D}_{{\bm{x}}}^{1},\ldots,\mathcal{D}_{{\bm{x}}}^{t}$, the TTA method produces a sequence of classifiers $p_{\theta_{1}},\ldots,p_{\theta_{t}}$, where each $\theta_{k}$ is adapted using $\mathcal{D}_{{\bm{x}}}^{k}$. This stands in contrast to the static source model $p_{\theta_{0}}$, which is trained once on labeled training data $\mathcal{D}_{\text{train}}$ and remains fixed during deployment. For simplicity, we refer to the source model as $p_{0}$ and the adapted models as $p_{1},\ldots,p_{t}$ henceforth.
@@ 3 Sequential Testing for TTA Monitoring
We now detail our approach to risk monitoring for test-time adaptation (TTA) methods using sequential testing. We begin by extending the risk monitoring framework of Podkopaev and Ramdas [30] to a deployment setting in which the model is continuously updated (§ 3.1). Next, inspired by Amoukou et al. [1], we derive a sequential test for the running test risk that does not require access to labels on the test data stream (§ 3.2)—a key innovation that enables rigorous statistical testing in TTA settings where test labels are unavailable. We then propose a concrete instantiation of our unsupervised test based on model uncertainty (§ 3.3) and online calibration of thresholds (§ 3.4). Finally, we describe techniques to enhance the detection power of the proposed tests (§ A.1). Our approach to monitoring risks in TTA is summarized in Algo. 1.
@@ 3.1 Risk Monitoring under Model Adaptation
Unlike in the static model setting considered by Podkopaev and Ramdas [30], we are interested in scenarios where a classifier is being continuously updated using a TTA method. Hence, we are interested in monitoring the risk not of a static source model $p_{0}$, but rather of a sequence of models $p_{1:t}$. To this end we define the hypotheses tested by our TTA risk tracker as:
@@
which also enjoys strong PFA control guarantees when using conjugate-mixture empirical Bernstein bounds [13]. Throughout the rest of this section, we abbreviate ${\mathbf{z}}_{k}^{(k)}$ as ${\mathbf{z}}_{k}$ and we shorten the sequence notation from ${\mathbf{z}}_{1},\ldots,{\mathbf{z}}_{t}$ to ${\mathbf{z}}_{1:t}$ to ease the notational burden.
@@ 3.2 Unsupervised Risk Monitoring
We defer the full proof to § B.2. A similar bound was proposed by Amoukou et al. [1], though with some key differences, which we discuss in detail in § A.1 and § 4. Importantly, the bound $B_{t}$ from Proposition 1 depends only on the test loss proxies and the source loss, meaning its corresponding lower-bound confidence sequence $L_{t}^{b}$ can be evaluated using a combination of unlabeled test data ($\mathcal{D}_{{\bm{x}}}^{k}$) and labeled source data ($\mathcal{D}_{\text{cal}}$). This makes it suitable for our final proposed unsupervised alarm:
@@ 3.3 Uncertainty as Loss Proxy
After introducing a general loss proxy ${\mathbf{u}}$ in the previous section, we now present a concrete instantiation based on model uncertainty. Specifically, we define the proxy function using the maximum class probability as $g({\mathbf{x}},p):=1-\max_{c}p({\mathbf{x}})_{c}$. We choose uncertainty, firstly, due to it being easy to implement: it requires no modifications to the underlying model and avoids the need for additional components, unlike alternative proxies based on model disagreement [33] or separate error estimators [1, 4]. Secondly, for $0\text{-}1$ loss this score approximates the conditional risk, up to calibration error:
@@
Although the conditional risk approximation improves when $p$ is well-calibrated, we do not need to assume the model’s uncertainty is well-calibrated under model adaptation [48] nor under distribution shift [28], as some previous work has required [16]. Returning to Assumption 1, uncertainty is a useful loss proxy when it separates high-loss and low-loss instances for a carefully chosen threshold $\lambda_{k}$—a task known as failure prediction [4, 50]. Failure prediction boils down to the ability to rank the test instances according to their loss values, which is a much weaker requirement in comparison to calibration [11]. Before demonstrating empirically in § 5 that using model uncertainty in Proposition 1 yields valid and tight lower bounds when monitoring TTA performance, we describe our threshold selection mechanism below.
@@ 3.4 Online Threshold Calibration
We now describe our procedure for selecting the loss and proxy thresholds used in the lower bound from Proposition 1. This step is critical for the effectiveness of our risk monitoring tool: poorly chosen thresholds can yield bounds that are either invalid (i.e., violate Assumption 1) or vacuous (i.e., excessively loose). Since our goal is to simultaneously minimize false positives and false negatives (cf. Assumption 1), we determine the loss threshold $\tau\in(0,C)$ and the source proxy threshold $\lambda_{0}\in(0,1)$ by maximizing the F1 score based on the source model’s proxy:
@@
where $\text{TP}=\sum_{i=1}^{N_{\text{cal}}}\mathbbm{1}[u_{0,i}>\lambda,z_{0,i}>\tau],\;\text{FN}=\sum_{i=1}^{N_{\text{cal}}}\mathbbm{1}[u_{0,i}\leq\lambda,z_{0,i}>\tau],\;\text{FP}=\sum_{i=1}^{N_{\text{cal}}}\mathbbm{1}[u_{0,i}>\lambda,z_{0,i}\leq\tau]$ and $u_{0,i}\sim{\mathbf{u}}_{0}$ and $z_{0,i}\sim{\mathbf{z}}_{0}$ are proxy and loss realizations of the source model $p_{0}$ on samples in $\mathcal{D}_{\text{cal}}$, respectively. Similarly, to select test proxy thresholds $\lambda_{1:t}$ we maximize F1 score while keeping the loss threshold $\hat{\tau}$ fixed: $\hat{\lambda}_{k}:=\operatorname*{arg\,max}_{\lambda}\text{F1}(\lambda,\hat{\tau};\mathcal{D}_{\text{cal}},p_{k}),$ where F1 is computed from proxy $u_{0}^{(k)}\sim{\mathbf{u}}_{0}^{(k)}$ and loss $z_{0}^{(k)}\sim{\mathbf{z}}_{0}^{(k)}$ realizations of the adapted model $p_{k}$ on the (same) calibration dataset $\mathcal{D}_{\text{cal}}$ (since no test labels are available). We emphasize that continuously adapting the proxy threshold is essential for preserving an effective bound $B_{t}$ under model adaptation. Using a static threshold throughout the test stream is insufficient, as many TTA methods can affect the scale of the observed uncertainties. For example, TENT [42] tends to reduce uncertainty over time due to its entropy minimization objective. Our full threshold selection procedure is summarized in Algo. 2.
@@ 4 Related Work
TTA [44, 20] aims to improve model performance under distribution shift by updating the model using unlabeled test data. Classic approaches include recomputing normalization statistics [36, 24], optimizing unsupervised objectives such as test entropy [42, 26, 27], energy [47], or pseudo-labels [17, 19], or adapting the last layer [15, 3, 35]. However, recent work has identified scenarios where TTA methods are ineffective [49, 35], and even harmful, degrading performance below that of the unadapted source model [3, 10, 46, 27, 6, 31, 43, 29]. While some studies propose heuristic indicators of TTA failure, such as high gradient norms [27], or estimate test accuracy directly [18, 31], there remains no principled framework for detecting risk violations of TTA methods with theoretical guarantees. Risk monitoring via sequential testing has been proposed by Podkopaev and Ramdas [30], though in a setting where test stream labels are available and the model remains static. Most relevant to our work is that of Amoukou et al. [1], who extend [30] to the test scenario without labels. We further build upon their framework by: (i) incorporating model adaptation (§ 3.1); (ii) deriving an unsupervised bound on the expected loss, rather than only a bound on the probability of high loss (§ 3.2); (iii) using model uncertainty as a proxy for loss instead of a separate error estimator (§ 3.3); (iv) proposing a simpler calibration method (§ 3.4) and showing it’s effectiveness for $0\text{-}1$ loss (§ 5.1); and (v) providing theoretical insights into why effective monitoring of continuous losses necessitates a change in the tested hypothesis (§ A.1). Also related is work by Bar et al. [2], where a sequential test for TTA methods based on betting martingales [32] is proposed. However, their test is designed to detect changes in predictive entropy, which may or may not lead to a degradation in performance—unlike our method, which directly tests for performance drops. We provide further related works in Appendix E.

### v2

@@ 1 Introduction
Test-time adaptation (TTA) [xiao2024beyond] has proven to be a powerful paradigm for prolonging the life of a model subjected to distribution shift. TTA methods adapt model parameters online, using only test batches of features and no labels. By leveraging unsupervised objectives such as test-time entropy [Wang2021TentFT] or pseudo-label losses [wang2022continual], these methods effectively ‘fine-tune’ model parameters on unlabeled test data. Despite their stark potential to retain high performance under a variety of distribution shifts [zhao2023pitfalls], TTA methods can suffer performance drops under severe shifts or after prolonged adaptation. The TTA literature has documented a range of persistent failure cases in which the model collapses entirely, resulting in near-zero accuracy [press2024entropy]. Alarmingly, these failures often occur silently and prohibit TTA methods from safe deployment in practice.
Timely detection of performance degradations—whether due to harmful distribution shifts or adaptation collapse—is thus crucial for safe deployment. At the same time, however, falsely flagging that a model should be taken offline and retrained can incur significant, avoidable costs given the size of modern predictive models. Recently, sequential testing has emerged as a promising statistical framework for monitoring model performance over time [podkopaev2021tracking]. When a predefined risk (or error) threshold is exceeded, the monitoring tool triggers an alarm. By employing time-uniform confidence sequences [howard2021time], such tools provide rigorous guarantees on the false alarm rate, thereby minimizing unnecessary retraining with high probability.
However, existing sequential risk monitors either require access to ground truth labels in production [podkopaev2021tracking] or do not account for model adaptation [jpmorgan2024sequential]. In this paper, we extend sequential testing to the challenging setting of TTA. This enables us to track the risk of a continuously evolving model without ever observing test labels. Our main contributions are as follows:
@@
  • In § 3.2 and § A.1, we extend prior work on unsupervised risk monitoring [jpmorgan2024sequential] to enable effective tracking of risks most commonly used in TTA, such as classification error.
@@ 2 Preliminaries
When averaged across samples, this squared loss corresponds to the Brier score, a strictly proper scoring rule [gneiting2007strictly]. Hence it captures not only the classifier’s error but also its calibration. Due to this connection, we refer to this second loss as the ‘Brier loss’ for short.
@@
To track how well a deployed model is performing, podkopaev2021tracking propose a risk monitoring framework based on sequential testing [ramdas2023game]. The performance of a model $p$, in the presence of a labeled test stream, is tracked using the following sequential test:
@@
To give the test anytime-valid properties (e.g. arbitrary stopping and restarting), podkopaev2021tracking rely on confidence sequences, which extend traditional confidence intervals to the sequential setting and offer time-uniform coverage guarantees [darling1967confidence, howard2021time]. A sequence of model losses on test data is used to construct an anytime-valid lower bound $L_{t}$ on the true running test risk $\bar{R}_{t}$:
@@
$\hat{U}_{0}=\hat{R}_{0}(p;\mathcal{D}_{\text{cal}})+w_{0}\>,\quad\hat{L}_{t}=\frac{1}{t}\sum_{k=1}^{t}\hat{R}_{k}(p;\mathcal{D}_{{\bm{x}}y}^{k})-w_{t}\>,$
@@
Figure 1: Alarm $\Phi_{t}$ is raised at $t_{\text{min}}$ as the lower bound $L_{t}$ on the running test risk $\bar{R}_{t}$ exceeds the upper bound $U_{0}$ on the source risk $R_{0}$.
which ensures that performance degradations are not incorrectly detected, thereby avoiding unnecessary (and potentially costly) interventions on the model. Notably, this guarantee requires only that the loss function is bounded $\ell\in[a,b]$. This minimal assumption makes the method broadly applicable, as it imposes no constraints on the data distributions, the predictive model, or the nature of distribution shift (beyond independence). To maintain minimal assumptions, it is necessary to rely on a conjugate-mixture empirical Bernstein bound [howard2021time] when constructing a lower confidence sequence for the test risk $L_{t}$ (see § B.1 for more details).
@@
In test-time adaptation [Wang2021TentFT], the model parameters $\theta$ are updated online as the model observes batches of unlabeled test features. Specifically, given a sequence of unlabeled test batches $\mathcal{D}_{{\bm{x}}}^{1},\ldots,\mathcal{D}_{{\bm{x}}}^{t}$, the TTA method produces a sequence of classifiers $p_{\theta_{1}},\ldots,p_{\theta_{t}}$, where each $\theta_{k}$ is adapted using $\mathcal{D}_{{\bm{x}}}^{k}$. This stands in contrast to the static source model $p_{\theta_{0}}$, which is trained once on labeled training data $\mathcal{D}_{\text{train}}$ and remains fixed during deployment. For simplicity, we refer to the source model as $p_{0}$ and the adapted models as $p_{1},\ldots,p_{t}$ henceforth.
@@ 3 Sequential Testing for TTA Monitoring
We now detail our approach to risk monitoring for test-time adaptation (TTA) methods using sequential testing. We begin by extending the risk monitoring framework of podkopaev2021tracking to a deployment setting in which the model is continuously updated (§ 3.1). Next, inspired by jpmorgan2024sequential, we derive a sequential test for the running test risk that does not require access to labels on the test data stream (§ 3.2)—a key innovation that enables rigorous statistical testing in TTA settings where test labels are unavailable. We then propose a concrete instantiation of our unsupervised test based on model uncertainty (§ 3.3) and online calibration of thresholds (§ 3.4). Finally, we describe techniques to enhance the detection power of the proposed tests (§ A.1). Our approach to monitoring risks in TTA is summarized in Algo. 1.
@@ 3.1 Risk Monitoring under Model Adaptation
Unlike in the static model setting considered by podkopaev2021tracking, we are interested in scenarios where a classifier is being continuously updated using a TTA method. Hence, we are interested in monitoring the risk not of a static source model $p_{0}$, but rather of a sequence of models $p_{1:t}$. To this end we define the hypotheses tested by our TTA risk tracker as:
@@
which also enjoys strong PFA control guarantees when using conjugate-mixture empirical Bernstein bounds [howard2021time]. Throughout the rest of this section, we abbreviate ${\mathbf{z}}_{k}^{(k)}$ as ${\mathbf{z}}_{k}$ and we shorten the sequence notation from ${\mathbf{z}}_{1},\ldots,{\mathbf{z}}_{t}$ to ${\mathbf{z}}_{1:t}$ to ease the notational burden.
@@ 3.2 Unsupervised Risk Monitoring
We defer the full proof to § B.2. A similar bound was proposed by jpmorgan2024sequential, though with some key differences, which we discuss in detail in § A.1 and § 4. Importantly, the bound $B_{t}$ from Proposition 1 depends only on the test loss proxies and the source loss, meaning its corresponding lower-bound confidence sequence $L_{t}^{b}$ can be evaluated using a combination of unlabeled test data ($\mathcal{D}_{{\bm{x}}}^{k}$) and labeled source data ($\mathcal{D}_{\text{cal}}$). This makes it suitable for our final proposed unsupervised alarm:
@@ 3.3 Uncertainty as Loss Proxy
After introducing a general loss proxy $\mathbf{u}$} in the previous section, we now present a concrete instantiation based on model uncertainty. Specifically, we define the proxy function using the maximum class probability as $g({\mathbf{x}},p):=1-\max_{c}p({\mathbf{x}})_{c}$. We choose uncertainty, firstly, due to it being easy to implement: it requires no modifications to the underlying model and avoids the need for additional components, unlike alternative proxies based on model disagreement [rosenfeld2023almost] or separate error estimators [jpmorgan2024sequential, corbiere2019addressing]. Secondly, for $0\text{-}1$ loss this score approximates the conditional risk, up to calibration error:
@@
Although the conditional risk approximation improves when $p$ is well-calibrated, we do not need to assume the model’s uncertainty is well-calibrated under model adaptation [zhang2024come] nor under distribution shift [ovadia2019can], as some previous work has required [kirsch2022a]. Returning to Assumption 1, uncertainty is a useful loss proxy when it separates high-loss and low-loss instances for a carefully chosen threshold $\lambda_{k}$—a task known as failure prediction [corbiere2019addressing, zhu2022rethinking]. Failure prediction boils down to the ability to rank the test instances according to their loss values, which is a much weaker requirement in comparison to calibration [guo2017calibration]. Before demonstrating empirically in § 5 that using model uncertainty in Proposition 1 yields valid and tight lower bounds when monitoring TTA performance, we describe our threshold selection mechanism below.
@@ 3.4 Online Threshold Calibration
We now describe our procedure for selecting the loss and proxy thresholds used in the lower bound from Proposition 1. This step is critical for the effectiveness of our risk monitoring tool: poorly chosen thresholds can yield bounds that are either invalid (i.e., violate Assumption 1) or vacuous (i.e., excessively loose). Since our goal is to simultaneously minimize false positives and false negatives (cf. Assumption 1), we determine the loss threshold $\tau\in(0,M)$ and the source proxy threshold $\lambda_{0}\in(0,1)$ by maximizing the F1 score based on the source model’s proxy:
@@
where $\text{TP}=\sum_{i=1}^{N_{\text{cal}}}\mathbbm{1}[u_{0,i}>\lambda,z_{0,i}>\tau],\;\text{FN}=\sum_{i=1}^{N_{\text{cal}}}\mathbbm{1}[u_{0,i}\leq\lambda,z_{0,i}>\tau],\;\text{FP}=\sum_{i=1}^{N_{\text{cal}}}\mathbbm{1}[u_{0,i}>\lambda,z_{0,i}\leq\tau]$ and $u_{0,i}\sim{\mathbf{u}}_{0}$ and $z_{0,i}\sim{\mathbf{z}}_{0}$ are proxy and loss realizations of the source model $p_{0}$ on samples in $\mathcal{D}_{\text{cal}}$, respectively. Similarly, to select test proxy thresholds $\lambda_{1:t}$ we maximize F1 score while keeping the loss threshold $\hat{\tau}$ fixed: $\hat{\lambda}_{k}:=\operatorname*{arg\,max}_{\lambda}\text{F1}(\lambda,\hat{\tau};\mathcal{D}_{\text{cal}},p_{k}),$ where F1 is computed from proxy $u_{0}^{(k)}\sim{\mathbf{u}}_{0}^{(k)}$ and loss $z_{0}^{(k)}\sim{\mathbf{z}}_{0}^{(k)}$ realizations of the adapted model $p_{k}$ on the (same) calibration dataset $\mathcal{D}_{\text{cal}}$ (since no test labels are available). We emphasize that continuously adapting the proxy threshold is essential for preserving an effective bound $B_{t}$ under model adaptation. Using a static threshold throughout the test stream is insufficient, as many TTA methods can affect the scale of the observed uncertainties. For example, TENT [Wang2021TentFT] tends to reduce uncertainty over time due to its entropy minimization objective. Our full threshold selection procedure is summarized in Algo. 2.
@@ 4 Related Work
TTA
[xiao2024beyond, liang2024comprehensive] aims to improve model performance under distribution shift by updating the model using unlabeled test data. Classic approaches include recomputing normalization statistics [schneider2020improving, nado2020evaluating], optimizing unsupervised objectives such as test entropy [Wang2021TentFT, niu2022efficient, niu2023towards], energy [yuan2024tea], or pseudo-labels [lee2013pseudo, liang2020we], or adapting the last layer [iwasawa2021test, boudiaf2022parameter, schirmer2024test]. However, recent work has identified scenarios where TTA methods are ineffective [zhao2023pitfalls, schirmer2024test], and even harmful, degrading performance below that of the unadapted source model [boudiaf2022parameter, gong2022note, yuan2023robust, niu2023towards, dobler2023robust, press2024entropy, wu2023uncovering, park2024medbn]. While some studies propose heuristic indicators of TTA failure, such as high gradient norms [niu2023towards], or estimate test accuracy directly [lee2024aetta, press2024entropy], there remains no principled framework for detecting risk violations of TTA methods with theoretical guarantees.
Risk monitoring via sequential testing
has been proposed by podkopaev2021tracking, though in a setting where test stream labels are available and the model remains static. Recent important extensions include [timans2025on], which proposes supervised risk monitoring for the more challenging setting of instantaneous risk control, and [prinster2025watch], which tackles the task using weighted-conformal martingales. Most relevant to our work is that of jpmorgan2024sequential, who extend [podkopaev2021tracking] to the test scenario without labels. We further build upon their framework by: (i) incorporating model adaptation (§ 3.1); (ii) deriving an unsupervised bound on the expected loss, rather than only a bound on the probability of high loss (§ 3.2); (iii) using model uncertainty as a proxy for loss instead of a separate error estimator (§ 3.3); (iv) proposing a simpler calibration method (§ 3.4) and showing it’s effectiveness for $0\text{-}1$ loss (§ 5.1); and (v) providing theoretical insights into why effective monitoring of continuous losses necessitates a change in the tested hypothesis (§ A.1). Also related is work by bar2024protected, where a sequential test for TTA methods based on betting martingales [ramdas2023game] is proposed. However, their test is designed to detect changes in predictive entropy, which may or may not lead to a degradation in performance—unlike our method, which directly tests for performance drops. We provide further related works in Appendix E.